An electric bulb is rated 220 V and 100 W.


When it is operated on 110 V, the power


consumed will be :-


Grade:

3 Answers

Narayanarao laveti
11 years ago

power (watts) = current (amps) x potential difference (volts)

P=I*V

=>P1/P2=V1/V2

=>220 v/110v=100w/v2

=>v2=100/2

:. V2=50 v

rohit yadav
11 years ago

25 W

Roshan Mohanty
11 years ago

P=V*I
I will be same in both cases
so
P1=V1*I
P2=V2*I
Dividing both equations,
P1/P2=V1/V2
100/P2=220/110
P2=50 (Ans)